OpenStudy (anonymous):

-3x+9y=9 Solve for y . Isn;t the answer y=9+1/3x ?

OpenStudy (cgreenwade2000):

Tell me how you got your answer.

OpenStudy (anonymous):

start by adding 3x to both sides what do you get

OpenStudy (anonymous):

Or you could start by dividing both sides by 3.

OpenStudy (anonymous):

@cgreenwade2000 -3x+9y=9 ; 9y=9+3x ; y = 9+1/3x .

OpenStudy (anonymous):

& Okay @moser90 & Okay @henpen

OpenStudy (cgreenwade2000):

@Isabel.A So you divided everything on both side by 9...... Except 9.

OpenStudy (anonymous):

so now you have 9y = 9 + 3x

OpenStudy (anonymous):

Ohhh Ohpse ! y=1+1/3x

OpenStudy (anonymous):

\[\text{Given}\\\ \ -3x+9y=9\\\text{Factor out 3 from the left-hand side.}\\\ \ \ 3(-x+3y)=9\\\text{Divide both sides by 3.}\\\ \ -x+3y=3\\\text{Add x to both sides to isolate 3y.}\\\ \ \ 3y=3+x\\\text{Divide both sides by 3 to isolate y.}\\\ \ \ y=\frac{3+x}3=\frac{x}3+1\]

OpenStudy (anonymous):

\[-3x+9y=9\]\[-x+3y=3\]\[3y=3+x\]\[y=1+x/3\]

OpenStudy (anonymous):

@moser90 Yeah .

OpenStudy (cgreenwade2000):

Helping someone with this problem doesn't take 100 people guys. Especially when it's clear to see she just missed a step.

OpenStudy (anonymous):

don't have to be rude she is responding to all of us so should we just ignore her

OpenStudy (anonymous):

._. Im confused @oldrin.bataku & Samee @henpen

OpenStudy (anonymous):

I Appericate ALL The Help .

OpenStudy (anonymous):

@cgreenwade2000 Now I Have y=1+1/3x Cause I Fixed it .

OpenStudy (anonymous):